OpenStudy (anonymous):

Help with equations! Thanks! TRANSLATE THE WORD PROBLEM BELOW INTO AN EQUATION; THEN SOLVE. Motor scooter #1 and motor scooter #2 are headed straight toward each other. Scooter #1 is traveling at 22 mph, and scooter #2 is traveling at 24 mph. If the two scooters are 92 miles apart, how many hours will it be before the meet?

OpenStudy (jziggy):

Since they're headed straight toward each other, we can easily figure out the rate at which the distance between them is decreasing. (22+24)mph = 46 mph Now you can easily set up an equation, 92 - 46x = 0 where x is measured in hours. This could also be expressed in the less simplified form, 92 - (22+24)x = 0